OpenStudy (anonymous):

(Checking to see if right) Its positive 3 right? Simplify negative 9 over 6 divided by 3 over negative 2.. 3 1 −1 −3

OpenStudy (anonymous):

question marks are negatives

OpenStudy (texaschic101):

(-9/6) / (-3/2) -- flip what you are dividing by, then multiply -9/6 * -2/3 = 18/18 which = 1